### v1

# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, skipabove=10pt, skipbelow=10pt

Abstract
Designing controllers that are both safe and performant is inherently challenging. This co-optimization can be formulated as a constrained optimal control problem, where the cost function represents the performance criterion and safety is specified as a constraint. While sampling-based methods, such as Model Predictive Path Integral (MPPI) control, have shown great promise in tackling complex optimal control problems, they often struggle to enforce safety constraints. To address this limitation, we propose DualGuard-MPPI, a novel framework for solving safety-constrained optimal control problems. Our approach integrates Hamilton-Jacobi reachability analysis within the MPPI sampling process to ensure that all generated samples are provably safe for the system. On the one hand, this integration allows DualGuard-MPPI to enforce strict safety constraints; at the same time, it facilitates a more effective exploration of the environment with the same number of samples, reducing the effective sampling variance and leading to better performance optimization. Through several simulations and hardware experiments, we demonstrate that the proposed approach achieves much higher performance compared to existing MPPI methods, without compromising safety.
I Introduction
Co-optimizing safety and performance is a critical challenge in the design of controllers for autonomous systems, especially in safety-critical domains such as autonomous vehicles, UAVs, and robotics. In such settings, controllers must ensure that performance objectives (such as efficiency, speed, or agility) are met while guaranteeing that safety constraints are never violated. Achieving this balance can be viewed as a constrained optimal control problem, where the goal is to satisfy safety constraints throughout the trajectory while optimizing performance objectives.
Several methods have been proposed to address this co-optimization challenge. Dynamic programming (DP) offers a rigorous solution to constrained optimization [1, 2]. Still, it is computationally intractable for many real-time applications due to the “curse of dimensionality”, often associated with DP-based solutions. On the other hand, MPC is more feasible for real-time applications, leveraging its ability to generate optimized control sequences over a receding horizon [3, 4]. Among these, Model Predictive Path Integral (MPPI) control is a sampling-based MPC method that has gained significant popularity due to its flexibility in handling complex dynamics, uncertainties, and non-linear systems [5, 6]. MPPI leverages stochastic sampling to optimize control trajectories, offering a scalable and efficient way to generate control actions. However, ensuring safety within MPPI has proven challenging, as the basic framework does not inherently account for hard safety constraints. Consequently, various extensions have been proposed to incorporate safety, each solving the co-optimization problem differently.
Classical MPPI-based approaches encourage safety constraints by penalizing unsafe sampled trajectories in the cost function, such that they are effectively ignored in the sample aggregation process. However, this approach cannot guarantee safety and its satisfaction depends on the penalty function. Another drawback is that computation is wasted on ignored samples, reducing the effective number of samples used for optimization. To enforce safety, approaches such as safety filtering and barrier functions have emerged, where constraints are enforced as a post-optimization correction [7, 8]. Such approaches solve the MPPI problem first and apply safety filtering afterward, or combine these two approaches [9, 10, 11].However, these approaches tend to ignore the long-term effect of safety action on the performance. Moreover, the overall performance of the system is still limited by the underlying MPPI algorithm. Other procedures use probabilistic methods to address the safety-performance co-optimization, such as modifying the sampling distribution to favor safer trajectories [12], using stochastic safety certificates to steer away from unsafe regions [13], or improving safety indirectly by enhancing adaptability and robustness to uncertainty [14, 15].
In this work, we introduce DualGuard-MPPI, a novel MPPI algorithm that incorporates safety constraints directly into the MPPI algorithm using Hamilton-Jacobi (HJ) reachability. Our key idea is to incorporate two safety filtering stages in the MPPI algorithm: first, we perform safety filtering along the sampled control perturbations, leading to provably safe “hallucinations” of the system trajectories, which are then used to optimize the performance objective. Second, we apply safety filtering on the selected control sequence before it is applied to the system to filter potential multi-modal sequences that may compromise safety when combined. The proposed approach imposes safety constraints throughout the trajectory optimization and control application stages, ensuring that the system operates within safe bounds at all times, without requiring any safety penalty tuning. Moreover, since all generated samples are provably safe, they all contribute to performance optimization, effectively reducing the MPPI sampling variance and leading to a better performance, as compared to vanilla MPPI algorithm.
In summary, the key contributions of this work are:
  • An MPPI framework that co-optimizes safety and performance while ensuring safety at all times;
  • Elimination of safety penalty tuning within MPPI and improvement in sampling efficiency by guaranteeing that all sampled control sequences are safe.
  • Extensive simulations and hardware experiments demonstrating the superior performance and robustness of the proposed approach without compromising safety.
II Problem Statement
Consider an autonomous system with state $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ that evolves according to dynamics $\dot{\mathbf{x}}=f(\mathbf{x},u,d)$, where $u\in\mathcal{U}$ and $d\in\mathcal{D}$ are the control and disturbance of the system, respectively. $d$ can represent potential model uncertainties or an actual, adversarial exogenous input to the system. We assume the dynamics are uniformly continuous in $u$ and $d$, bounded, and Lipschitz continuous in $\mathbf{x}$ for fixed $u$ and $d$. Finally, let $\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}(\tau)$ denote the system state at time $\tau$, starting from the state $\mathbf{x}$ at time $t$ under control signal $\mathbf{u}(\cdot)$ and disturbance signal $\mathbf{d}(\cdot)$ while following the dynamics. A control signal $\mathbf{u}(\cdot)$ is defined as a measurable function mapping from the time horizon to the set of admissible controls $\mathcal{U}$, and a disturbance signal is similarly defined. We additionally assume that the control and disturbance signals $\mathbf{u}(\cdot)$ and $\mathbf{d}(\cdot)$ are piecewise continuous in $t$. This assumption ensures that the system trajectory $\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}}$ exists and is unique and continuous for all initial states [16, 17].
In addition, we are given a failure set $\mathcal{F}\subset X$ that the system must avoid at all times (e.g., obstacles for a navigation robot). The safety constraint is encoded via a Lipschitz continuous function $l(\cdot):\mathcal{F}=\{\mathbf{x}:l(\mathbf{x})\leq 0\}$. We aim to design a controller that optimally balances the system’s safety constraints and performance objectives. The performance objective is given by minimizing a cost function $S$ over the system trajectory, given by:
$S(\xi)=\phi(\mathbf{x}(t_{f}),t_{f})+\int_{0}^{t_{f}}{\mathcal{L}}(\mathbf{x}(t),u(t),t)\mathrm{d}t$ (1)
where $\phi$ and $\mathcal{L}$ represent the terminal and running cost respectively, and $t_{f}$ the task completion time. Specifically, we aim to ensure that the control actions $u$ drive the system towards minimizing $S$ while rigorously maintaining system safety by preventing the state $\mathbf{x}$ from entering $\mathcal{F}$ even for the worst case disturbances $d$. This takes the form of the following constrained optimal control problem:
$u^{*}(\cdot)$ $=\operatornamewithlimits{argmin}_{u(\cdot)}S(\xi_{\mathbf{x},t}^{\mathbf{u},\mathbf{d}},u(\cdot))$ (2)
subject to:
$\dot{\mathbf{x}}(t)$ $=f(\mathbf{x}(t),u(t),d(t)),$
$\quad\;l(\mathbf{x}(t))$ $>0,\;u(t)\in\mathcal{U},\;d(t)\in\mathcal{D},\forall t\in[t,t_{f}]$
the optimization problem defined in (2) in general is non-convex and can be difficult to solve. In this work, we propose a novel MPPI method to solve this problem.
III Background
III-A Model Predictive Path Integral (MPPI)
MPPI [5, 6] aims to solve a classical stochastic optimal control problem using a zeroth-order sampling-based MPC scheme. Specifically, at each state, MPPI samples K random control sequences around a nominal control sequence for a time horizon $\hat{T}\leq t_{f}$ discretized over $H$ evenly spaced steps. The control input for the kth sequence at jth time step is given by: $u^{k}(x,j)=u(x,j)+\delta_{j}^{k}$, where $u(x,j)$ is a nominal control sequence and $\delta_{j}^{k}$ a randomly sampled control perturbation. Next, the state trajectory $\xi_{j}^{k}$ and the cost to go $S(\xi_{j}^{k})$ corresponding to each control sequence is computed. It is shown that the optimal control sequence can be approximated by the following update law [9]:
$u(\mathbf{x},j)^{\ast}\approx u(\mathbf{x},j)+{\frac{\sum_{k=1}^{K}\exp[-(1/\lambda)S(\xi_{j}^{k})]\delta_{j}^{k}}{\sum_{k=1}^{K}\exp[-(1/\lambda)S(\xi_{j}^{k})]}}$ (3)
where $\lambda\in\mathbb{R}^{+}$ is the temperature parameter that weighs the contributions from different control sequences. The first control in the optimal sequence is applied to the system and the entire process is repeated at the next time step.
{mdframed}
[style=MyFrame,nobreak=false] Running example (Safe Planar Navigation): As a running example we consider a Dubins’ car trying to reach a goal in a cluttered environment. The system can be represented by the following three dimensional dynamics:
$\dot{\mathbf{x}}=\begin{bmatrix}\dot{x}&\dot{y}&\dot{\theta}\end{bmatrix}=\begin{bmatrix}V\cos(\theta)&V\sin(\theta)&u\end{bmatrix}$ (4)
[…]
Dotted lines in Fig. 1 represent possible rollouts of the system following a randomly perturbed control sequence starting from the current state. We refer to these as the hallucinations of the system for a given state.
Figure 1: Visualization of hallucination step in the MPPI algorithm.
The hallucinations that fail to avoid obstacles achieve a higher cost (shown in red). Thus, the update law in (3) prioritizes control perturbations that will keep the car safe (shown in a blue-to-purple gradient, with blue indicating lower costs). However, in general, it could be possible that all samples lead to safety violations. Moreover, if the number of safe samples is small, it could lead to a high variance in the performance of MPPI, necessitating the need to account for safety constraints rigorously.
III-B Hamilton-Jacobi Reachability
One way to guarantee safe operation of autonomous continuous-time dynamical systems is through Hamilton-Jacobi (HJ) reachability analysis. This approach involves computing the Backward Reachable Tube (BRT) of the failure set $\mathcal{F}$. The BRT captures the states from which the system is not able to avoid entering $\mathcal{F}$ within some time horizon $T$, despite the best control effort.
In HJ reachability, the BRT computation is formulated as a zero-sum game between control and disturbance. Specifically, a cost function is defined to capture the minimum distance to $\mathcal{F}$ over time:
$J(\mathbf{x},t,u(\cdot),d(\cdot))=\min_{\tau\in[t,T]}l(\mathbf{x}(\tau))$ (5)
The goal is to capture this minimum distance for optimal system trajectories. Thus, we compute the optimal control that maximizes this distance (drives the system away from the failure set) and the worst-case disturbance signal that minimizes the distance. The value function corresponding to this robust optimal control problem is:
$V(\mathbf{x},t)=\adjustlimits{\inf}_{\gamma\in\Gamma(t)}{\sup}_{u(\cdot)}\{J(\mathbf{x},t,u(\cdot),\gamma[u](\cdot))\},$ (6)
where $\Gamma(t)$ defines the set of non-anticipative strategies for the disturbance [18]. The value function in (6) can be computed using dynamic programming, which results in the following final value Hamilton-Jacobi-Isaacs Variational Inequality (HJI-VI) [18, 19, 20]:
[…]
$V(\mathbf{x},T)=l(\mathbf{x}),\quad\text{for}\ t\in\left[0,T\right]$
$D_{t}$ and $\nabla$ represent the time and spatial gradients of the value function. $H$ is the Hamiltonian, which optimizes over the inner product between the spatial gradients of the value function and the dynamics:
$H(\mathbf{x},t,\nabla V(\mathbf{x},t))=\min_{u\in\mathcal{U}}\max_{d\in\mathcal{D}}\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d)$ (8)
Once the value function is obtained, the BRT is given as the sub-zero level set of the value function:
$\mathcal{V}(t)=\{\mathbf{x}:V(\mathbf{x},t)\leq 0\}$ (9)
The corresponding optimal safe control can be derived as:
$u^{*}_{safe}(\mathbf{x},t)=\operatornamewithlimits{argmax}_{u\in\mathcal{U}}\min_{d\in\mathcal{D}}\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d)$ (10)
In safety-ensuring applications, we want to guarantee that the system does not enter $\mathcal{F}$ at any time; for this reason, we use a time-converged BRT, as the set of unsafe states often stops growing after some amount of time. Consequently, we can use the converged value function $V(\mathbf{x})$, which, when used to synthesize safety controllers, gives an expression identical to (10) without its time dependency.
To compute the value function and obtain the BRT and the optimal controller, we can rely on methods that solve the HJI-VI in (6) numerically [21, 22] or using learning-based methods [23, 24].
III-C Least Restrictive Filtering
If the BRT and associated converged value function are known, a Least Restrictive Filter (LRF) can be deployed to guarantee the safe operation of the system. The LRF is constructed as follows:
$u(\mathbf{x},t)=\begin{cases}u_{\text{nom}}(\mathbf{x},t)&V(\mathbf{x})>0\\ u^{*}_{\text{safe}}(\mathbf{x})&V(\mathbf{x})=0\end{cases}$ (11)
Here, $u_{\text{nom}}(\mathbf{x},t)$ corresponds to an arbitrary nominal controller that might optimize a performance criterion without enforcing safety constraints. This control is used when the value function $V(\mathbf{x})$ is positive, as the system is not at risk of breaching safety. Whenever the system reaches the boundary of the BRT ($V(\mathbf{x})=0$), it switches to $u^{*}_{safe}(x)$ (10) as this optimal control is guaranteed to maintain or increase $V(\mathbf{x})$ keeping the system from entering the unsafe states determined by the BRT, thereby enforcing safety at all times. We refer the reader to [7] for details on this filtering technique and proof of the safety guarantees.
{mdframed}
[style=MyFrame,nobreak=false]
Running example (Safe Planar Navigation):
We can use the HJI-VI in (6) to calculate the value function and associated BRT, over which we can implement a LRF to guarantee the system’s safety. In this case considering the dynamics in (4) and the environment definition, we compute the value function using the LevelSetToolbox[21].
In Fig. 2, the purple trajectory represents the system’s response to a perturbed control sequence, reaching a critical state at the boundary of the safe set (yellow dot). The boundary of the safe set at this state is illustrated with a dotted line. The red trajectory shows the continuation of the system using the perturbed controls, which leads to the failure set. By contrast, the blue trajectory demonstrates how the unsafe execution can be made safe. This is achieved by applying a LRF step to each control action in the sequence before execution.
Figure 2: Trajectory correction by LRF. System under perturbed controls (purple) reaches the safe set boundary (yellow dot). The red path shows an unsafe execution, while the blue path demonstrates how safety filtering maintains the system outside the failure set.
IV Dualguard MPPI
As discussed in Sec. III-A, MPPI solves the optimal control problem in (2) using a sampling-based method. While MPPI may encourage the satisfaction of safety constraints in (2) via penalizing safety violations in the cost function, ensuring safety constraints remains challenging. Moreover, by construction, the high-cost safety breaching sequences are mostly ignored during the optimal control sequence computation, wasting computational resources that could have been used to refine the system performance further. While the safety filtering mechanism discussed in III-C can provide a safety layer after MPPI to enforce the safety constraint, this approach is myopic in nature and might lead to performance impairment in favor of safety.
To overcome these challenges, we propose DualGuard MPPI, a two-layered safety filtering approach. First, the safety filtering is incorporated during the sampling process itself, where a least restrictive filter is applied along the sampled control sequence rollouts using a pre-computed safety value function. This ensures that all hallucinations satisfy the safety constraint and can contribute to performance optimization. To ensure that the resultant optimal control sequence also satisfies the safety constraint, we filter the output optimal control sequence by a safety filter as well. In addition to ensuring the safety constraints, the proposed framework leads to an increased sample efficiency as the safe hallucinations keep all sampled trajectories safe and relevant to the performance objective, thereby avoiding ”sample wastage” due to safety constraint violation. The proposed algorithm is presented in Alg. 1. Details on the proposed filtering stages are discussed next.
IV-A Generating Safe hallucinations
As in classical MPPI, we consider $K$ sequences of random perturbations $\delta_{j}^{k}$, that modify a nominal control sequence $u_{j}$. The perturbed sequences are applied to the dynamic model of the system while being filtered at each time step along the horizon, using the LRF in eq. (11), resulting in hallucinated trajectories that are guaranteed to maintain safety. The cost-to-go $S^{k}$ for the filtered trajectory is calculated over the safety-filtered control perturbation sequence $\Delta_{j}^{k}$.
The nominal control sequence $u_{j}$, the filtered control perturbations $\Delta_{j}^{k}$ and the cost over the safe hallucinated trajectories $S^{k}$ are used to calculate the optimal control sequence $u^{*}_{j}$ using the update rule defined in (3). As we weigh over controls that only produce safe trajectories, all K sequences contribute to the performance optimization, reducing the variance of MPPI algorithm and leading to better performance.
IV-B Output Least restrictive filtering
Even though the optimal control sequence $u^{*}_{j}$ was obtained by weighing control perturbations that individually resulted in safe trajectories, the safety guarantees may not hold for $u_{j}^{*}$. For this reason, before applying the first control in the sequence $u^{*}_{0}$ to the system, we perform one last LRF step to guarantee the safe operation of the complete system. The rest of the sequence is used as the nominal control for the next time step, and the entire process is repeated.
One case that encourages this last filtering stage corresponds to scenarios where the safe hallucinations present multimodality. We can imagine a vehicle trying to swerve around an obstacle by turning right or left; even if both swerving maneuvers result in safe behavior, a weighting over hallucinations split between these two modalities could result in maintaining a straight trajectory, causing a collision. Such issues are resolved by this additional filtering stage which picks one of the safe modes.
{mdframed}
[style=MyFrame,nobreak=false]
Running example (Safe Planar Navigation):
Considering our running example we observe the proposed steps of DualGuard MPPI in Fig. 3. First, we visualize the unmodified hallucination step in the left panel. Next, using the same control perturbations, we show how the safe hallucination step renders all the samples safe. Hallucinated trajectories are presented in a blue-to-red scale, corresponding to the associated low-to-high costs.
The right panel of Fig. 3 shows how the safe hallucinations might present themselves in a multimodal fashion for some obstacle configurations; this indicates a possible failure mode where the control sequence given by the update rule in (3) could drive the vehicle directly into the obstacles, leading to safety violations. The proposed output filtering stage safeguards against this possibility.
Figure 3: Unmodified hallucinations (Left). Safe hallucinations (Center). Possible multimodality on safe hallucinations encourages the use of output least restrictive filtering (Right).
V Simulation Studies
We now evaluate the performance of our proposed method across several simulation studies and on a real hardware testbed. For this, we compare the proposed approach against five baseline methods. Each baseline builds on the MPPI control approach but incorporates distinct modifications to enhance safety.
Obstacle Penalty (Obs. penalty): A classical MPPI approach to collision avoidance by adding a high penalty for entering the obstacles in the cost function, incentivizing the system to avoid these penalized regions.
BRT Penalty (BRT penalty): Advanced MPPI methods penalize hallucinated trajectories that enter a safe zone defined as augmented obstacles, Discrete Barrier States (DBaS), or Control Barrier Functions (CBF) [25, 26, 27, 28]. To contrast against this technique, we implement a baseline that penalizes hallucinated trajectories that enter the BRT of the obstacles. This approach is more preemptive, as it penalizes samples that have entered states from which such a collision is unavoidable but may have not yet collided with obstacles within the sampling horizon.
Obstacle Penalty with Output LRF (Obs. pen. + LRF): Extension of the Obstacle Penalty baseline by adding a least restrictive filtering step. This corresponds to a naive safety filtering approach where the least restrictive filter is used to correct any control that would result in a safety violation just before it is applied.
BRT Penalty with Output LRF (BRT pen. + LRF): Similar to the previous baseline, this extends the BRT Penalty baseline by adding least restrictive filtering safety guarantees.
Shield Model Predictive Path Integral (Shield MPPI): This method leverages the approach described in Shield-MPPI[10], where a cost related to the CBF condition is considered during hallucinations, followed by an approximate CBF-based repair step over the resulting updated control to promote safety. To ensure a fair comparison, we will use the HJ value function as a barrier function so that all methods work on the maximal safe set [29].
Evaluation Metrics: To quantitatively assess the performance of each method, we define the following evaluation metrics. These metrics will be measured over batches of simulation runs for each method, providing insights into safety, efficiency, and computational demands.
  • Failure: Percentage of episodes that result in a safety violation (e.g., a collision).
  • RelCost: Average normalized cost of the trajectory, using the proposed method as the baseline. We only consider successful executions across all methods for RelCost computation. This allows for a fair comparison and ensures that the RelCost is not affected by the high failure penalties for some of the baselines.
Furthermore, when deemed relevant, we define additional performance-related metrics specific to each case study.
V-A Safe Planar Navigation
We assess the running-example safe planar navigation with dynamics defined in (4). In each episode, the system must navigate from an initial state $\mathbf{x}_{0}$ to a goal region of radius 0.1 m in the $xy$ plane centered at the goal state $\mathbf{x}_{g}$, within a time horizon of $T=20\text{ s}$. The optimal control problem’s cost function is defined as the squared distance from the goal, a control effort penalty, and a safety penalty $P(\mathbf{x})$:
$S=\sum_{t=0}^{T}(\mathbf{x}_{t}-\mathbf{x}_{g})^{T}Q(\mathbf{x}_{t}-\mathbf{x}_{g})+m(u_{t})^{2}+P(\mathbf{x}_{t})\$ (12)
with $Q=\text{diag}([1,1,0])$, $m=0.2$, and $P=10000$ if the state is inside the obstacles or the BRT (corresponding to that controller’s cost penalty type) and $K=0$ otherwise.
We test each controller in the same set of 100 episodes, where each episode is defined by a particular pair $(\mathbf{x}_{0},\mathbf{x}_{g})$. These initial-goal state pairs were selected randomly but constrained to be outside the BRT, near the environment boundary, and separated by at least 5 m for diversity. The results are summarized in Table I. In addition to the metrics previously defined, we also compute the Success rate, defined as the percentage of episodes that reach the goal region safely (i.e., without collisions) within the time horizon, and Timeout, defined as the percentage of episodes that fail to reach the target within the given time horizon.
Figure 4: Safe Planar Navigation environment, with obstacles in gray and a slice of the BRT for a southwestern heading is in teal.
TABLE I: Safe Planar Navigation results over 100 episodes.
K Method Success Timeout Failure RelCost
1000 Obs. penalty 49 1 50 1.34 ± 1.19
BRT penalty 72 21 7 1.89 ± 2.11
Obs. pen. + LRF 80 20 0 1.35 ± 1.20
BRT pen. + LRF 74 26 0 1.90 ± 2.12
Shield MPPI 83 17 0 1.32 ± 0.87
DualGuard (Ours) 99 1 0 1.00
250 Obs. penalty 31 0 69 1.10 ± 0.24
BRT penalty 58 15 27 2.35 ± 3.40
Obs. pen. + LRF 81 19 0 1.27 ± 0.92
BRT pen. + LRF 69 31 0 1.86 ± 2.03
Shield MPPI 78 21 1 1.40 ± 0.86
DualGuard (Ours) 98 2 0 1.00
60 Obs. penalty 11 0 89 0.99 ± 0.03
BRT penalty 43 21 36 3.16 ± 4.17
Obs. pen. + LRF 70 30 0 0.99 ± 0.03
BRT pen. + LRF 55 45 0 3.16 ± 4.17
Shield MPPI 64 36 0 1.08 ± 0.14
DualGuard (Ours) 96 4 0 1.00
First examining the results for $K=1000$, we see that our proposed method achieves the lowest failure rate and highest success rate. While all methods with an output Least Restrictive Filter avoid collisions as expected, our method also times-out in fewer episodes, indicating a better performance. The safety-performance co-optimization capabilities of the proposed method are also evident from its lowest RelCost.
To understand the different controllers’ behavior, consider the trajectories shown in Figure 4. Here, the obstacle penalty method fails early by entering a region in the northeast corner which is impossible to escape and hence avoided by all BRT-aware methods. The other baseline methods execute a reasonable trajectory to the goal, but the proposed method manages to find a lower-cost trajectory by taking a shortcut south after rounding the first obstacle. This is because the baseline methods are still fundamentally limited by the exploration capabilities of MPPI – if the underlying samples are unsafe, they contribute little to the performance optimization. On the other hand, the proposed method generates safe hallucinations, allowing it to explore more relevant parts of the environment with the same number of samples, as previously visualized in Fig. 3, leading to a better performance optimization.
A consequence of the superior exploration capabilities of the proposed method is that it can achieve similar success rates with far fewer samples. We summarize the results for different methods as we vary the number of samples in Table I. We note that while all methods degrade with fewer samples as expected, DualGuard remains relatively consistent. With only 60 samples, it succeeds in 96% of the episodes, with the closest contender not going over 70%. This could enable the deployment of the proposed method on resource-constrained systems, which MPPI methods typically struggle with (as also indicated by the poor performance of other MPPI baselines with fewer samples).
V-B Autonomous Quadrotor Navigation
Figure 5: Quadrotor environment with obstacles (gray), initial position (olive), and goal (dark green). Trajectories for each method are shown in different colors, obstacle intrusions are denoted by red.
Next, we demonstrate our method on a 6-dimensional nonlinear quadrotor system. The system obeys the following dynamics:
$\dot{\mathbf{x}}$ $=\begin{bmatrix}\dot{x}&\dot{y}&\dot{z}&\dot{v_{x}}&\dot{v_{y}}&\dot{v_{z}}\end{bmatrix}^{T}$ (13)
$=\begin{bmatrix}v_{x}&v_{y}&v_{z}&a_{g}\tan u_{\theta}&-a_{g}\tan u_{\phi}&u_{T}-a_{g}\end{bmatrix}^{T}$
where $a_{g}=9.81\text{ m/s}^{2}$ is the acceleration due to gravity, and acceleration $u_{T}$, roll $u_{\phi}$, and pitch $u_{\theta}$ are the controls $u=[u_{T},u_{\phi},u_{\theta}]$ of the system with ranges $u_{T}\in[7.81,11.81]\text{ m/s}^{2}$ and $u_{\phi},u_{\theta}\in[-17.18^{\circ},17.18^{\circ}]$. Zero mean Gaussian noise is added to the actions to simulate real-world actuation errors.
The environment is a $10\text{m}\times 10\text{m}\times 4\text{m}$ room containing 10 spherical obstacles of varying sizes. The BRT of the system is computed using the LevelSetToolbox [21]. The task of the quadrotor is to reach a goal sphere with radius $0.3m$ around the goal state $\mathbf{x}_{g}$ (shown by the green sphere in Fig. 5), starting from a random initial position $\mathbf{x}_{0}$ while avoiding collisions with the obstacles. The system must reach the goal location within a time horizon of $T=10\text{s}$; otherwise, the episode is considered timeout. If the system collides with an obstacle or the environment boundary, the episode is considered unsafe.
The cost function for the MPPI algorithm consists of the distance to the goal location, control effort penalty, and a safety penalty $P(\mathbf{x})$:
$S=$ $(\mathbf{x}_{t}-\mathbf{x}_{g})^{T}Q(\mathbf{x}_{t}-\mathbf{x}_{g})$ (14)
$+\|(1/u_{range})^{T}(u_{t}-u_{mean})\|_{2}^{2}+P(\mathbf{x}_{t})$
where $Q=\text{diag}([20,20,20,0,0,0])$, $u_{range}=[2,0.3,0.3]$ and $u_{mean}=[9.81,0,0]$. $P(\mathbf{x})$ is 10 if the state is inside an obstacle or the BRT, corresponding to that controller’s cost penalty type. For all methods, $500$ hallucinations are simulated for a horizon of $0.3s$. Each method is tested starting from the same 100 randomly sampled initial positions over the environment with at least $5m$ distance from the goal location. The results are summarized in Table II.
In this case, the obstacle penalty method leads to a very poor success rate. It is partially due to the added disturbance in the system that makes it challenging to ensure safety. A BRT penalty improves the safety rate from 14% to 30%, but it still does not enforce safety, as shown in Fig. 5. By penalizing a decrease in safety in the cost function and further optimizing the MPPI solution to encourage safety, Shield MPPI decreases the number of collisions to 21. In contrast, safety is guaranteed by a LRF of the controls, resulting in a 100% safety rate. Nevertheless, a significant portion of the hallucinations end up inside obstacles and are unable to contribute to the control synthesis meaningfully. Therefore, methods with safety-filtered actions experience many trajectories that fail to reach the goal within the given task horizon despite maintaining safety, in Fig. 5 we show such timeout executions. By filtering the hallucinations, our method incorporates the rollouts that would otherwise be unmeaningful into the optimization. Doing so achieves the best success and safety rates among all methods while achieving a similar cost on successful trajectories compared to the others that guarantee safety.
TABLE II: Quadrotor simulation results.
Method Success Timeout Failure RelCost
Obs. penalty 14 0 86 0.43 $\pm$ 0.05
BRT penalty 30 0 70 0.50 $\pm$ 0.07
Obs. pen. + LRF 68 32 0 1.08 $\pm$ 0.35
BRT pen. + LRF 65 35 0 1.02 $\pm$ 0.38
Shield MPPI 69 10 21 0.84 $\pm$ 0.18
DualGuard (Ours) 75 25 0 1.00
VI Hardware Experiments - RC Car
[…]
${\dot{\mathbf{x}}=\begin{bmatrix}\dot{x}&\dot{y}&\dot{\theta}\end{bmatrix}=\begin{bmatrix}V\cos(\theta)+d_{x},&V\sin(\theta)+d_{y},&V\tan(\delta)/L\end{bmatrix}}$ (15)
We adapt our evaluation metrics for this hardware study to better reflect real-world, single-run applicability. Instead of batch statistics, we measure the CompTime for each method – the time taken to generate and evaluate potential samples within a multiple-lap run. This metric reflects how well-suited each technique is for real-time control. We also report the car’s average Speed over three laps to measure how aggressive the policy is. The RelCost metric remains consistent with the simulations, providing a normalized cost relative to the proposed approach for the methods that managed to maintain safety.
[…]
$\small S=(V_{max}-V)^{2}+K_{c}(l_{center}-l(x))+P(\mathbf{x})$ (16)
The controllers were implemented using JAX [30] on a laptop equipped with an NVIDIA GeForce RTX 4060. We generate $1000$ parallel hallucinations (with $100$ time steps each) in a loop running at $50Hz$. Results are summarized in Table III, and trajectories for the first lap are shown in Fig 6.
First, we highlight the need for hard safety constraints as the methods that only rely on safety penalties fail to clear the top-left tight turn in the track as shown in Fig 6. Fine-tuning the cost function and MPPI parameters might allow unfiltered methods to complete laps. Still, we want to consider and compare methods that provably allow for safe executions.
TABLE III: Hardware experiments results summary.
Method CompTime (ms) RelCost Speed (m/s)
Obs costs 1.8 ± 0.3 fail 1.00 ± 0.05
BRT costs 1.8 ± 0.3 fail 1.01 ± 0.06
Obs costs + LRF 1.7 ± 0.4 1.1874 1.03 ± 0.12
BRT costs + LRF 1.8 ± 0.4 1.1626 1.04 ± 0.12
Shield-MPPI 1.7 ± 0.2 1.1038 1.04 ± 0.08
DualGuard (Ours) 2.5 ± 0.4 1.0000 1.10 ± 0.11
The proposed method leads to faster and more performant trajectories than the other safe baselines. A direct comparison with the baselines that also use an output LRF illustrates that the proposed safe hallucination step improved the quality of the samples as exemplified in Fig. LABEL:fig:exp_main_result(B)(C), leading to a better overall performance and a higher average speed. Also, the proposed method outperforms the Shield-MPPI baseline even after tuning its hyperparameters to the best of our capabilities so that it maintains safety without an excessive impact on performance.
The computation times are nearly identical across all baselines, as each method fundamentally involves calculating performant terms of the cost function and querying the obstacle set or BRT for safety-related penalties. The proposed method introduces an additional LRF step for each sample along hallucinated trajectories, resulting in a slight increase in computational time. Nevertheless, all methods, including the proposed one, operate well within the $20ms$ time budget, leaving ample time for the control loop to handle state estimation, communications, and actuation.
Figure 6: Top view of the RC car’s trajectories under each method.
VII Limitations and Future Work
In this work, we have presented DualGuard-MPPI, a novel MPPI framework for addressing the safety-constrained optimal control problem. By integrating Hamilton-Jacobi reachability analysis with MPPI, our approach ensures strict adherence to safety constraints throughout both the sampling and control execution phases. This combination enables high-performance trajectory optimization without compromising safety, validated through extensive simulation and real-world experiments. DualGuard-MPPI stands out for its capability to eliminate safety-related terms from the cost function, thereby streamlining the optimization process to focus purely on performance objectives.
While these contributions establish DualGuard-MPPI as a robust and scalable framework, certain limitations remain that warrant further exploration. The requirement for a precomputed BRT introduces considerable computational overhead during setup, which may hinder its implementability in rapidly changing environments. Additionally, the reliance on explicitly defined system dynamics may restrict the framework’s applicability to systems with highly complex or partially unknown models. Addressing these challenges—such as by leveraging online reachability methods to dynamically update BRTs [31, 32, 33] or using reachability methods for black-box systems [34, 35] could significantly expand the framework’s usability. In addition, we will explore the deployment of the proposed approach on other safety-critical robotics applications.